OpenStudy (anonymous):

Solve for X: 2/3(x-4)=2x I got -2 but I don't think I am right.

OpenStudy (popitree):

\[\frac{ 2 }{ 3(x-4) } = 2x\]

OpenStudy (popitree):

is this the equation?

OpenStudy (anonymous):

No its like 2 over 3 * (x-4) =2x

jhonyy9 (jhonyy9):

2(x-4) ------ = 2x 3 this wann be there

OpenStudy (anonymous):

No its like 2/3 multiplied by (x-4) = 2x

jhonyy9 (jhonyy9):

2(x-4) 2x ------ = ---- 3 1 cross multiplie and so what will get ?

OpenStudy (anonymous):

I ended up with negative 2 but I do not think I am correct.

jhonyy9 (jhonyy9):

how you have got it ?

OpenStudy (anonymous):

So we would put 3/2 and multiply it by 2x/1

jhonyy9 (jhonyy9):

2(x-4) 2x ------ = ---- 3 1 cross multiplie what will get ?

jhonyy9 (jhonyy9):

1*2(x-4) = 3*2x yes ?

OpenStudy (anonymous):

Okay... just got 1 but thats definitely not the answer....

jhonyy9 (jhonyy9):

please try it with me now step by step

OpenStudy (anonymous):

Okay so then we would subtract so -2 is right :D

OpenStudy (anonymous):

1*2(x-4)=3*2x 2x-8=6x -2 -2 -8/4=4x/4 -2 = x :D

OpenStudy (anonymous):

Is that right?

jhonyy9 (jhonyy9):

x=-2 is right sure

OpenStudy (anonymous):

Thank you :)!!! <3

jhonyy9 (jhonyy9):

np was my pleasure